The perimeter of reqular pentagon is 22 centimeters. How long is one side?
Found 2 solutions by praseenakos@yahoo.com, rajagopalan:
Answer by praseenakos@yahoo.com(507)

QUESTION

The perimeter of reqular pentagon is 22 centimeters. How long is one side?


ANSWER:

A pentagon has five sides.

All the five sides are of equal length for a REGULAR PENTAGON.

Perimeter = sum of its five sides.


Since all sides are equal....let us assume that each side is 'x'centimeters.
then perimeter = x + x + x + x + x = 5x


It is given that perimeter = 22 centimeters.


==> 5x = 22


Divide both sides by 5


==> 5x/5 = 22/5



==> x = 4.4 cm


So each side is of 4.4 cm long.


Hope you found the explanation useful....


Regards.

Praseena.

Answer by rajagopalan(174)

The perimeter of reqular pentagon is 22 centimeters.
length of one side= perimeter/number of sides(5)= 22/5=4.4 cm.
OKAY